OpenStudy (chosenmatt):

Which is 625 written as a power of a number? A. 34 B. 35 C. 53 D. 54

OpenStudy (phi):

if you have exponents you should use ^ to show them Example: the 34 is probably 3^4 , right ? it means 3*3*3*3 (3 times itself 4 times) if you work that out, you will see it is not 625

OpenStudy (anonymous):

A. 34

OpenStudy (chosenmatt):

thanks @phi

OpenStudy (anonymous):

maybe . . .

OpenStudy (cwrw238):

no itc ant be 3^4 3^4 = 3*3*3*3= 81

OpenStudy (cwrw238):

because the number ends in 5 it must be 5 to some power

OpenStudy (chosenmatt):

OKAY NOW IM CONFUSE @phi

OpenStudy (cwrw238):

well its either 5^3 or 5*4

OpenStudy (cwrw238):

work both of those out and you'll get the correct answer

OpenStudy (cwrw238):

find which one comes to 625

OpenStudy (cwrw238):

* tahts either 5^3 or 5^4

OpenStudy (chosenmatt):

it's 5^4

OpenStudy (cwrw238):

yes 5*5*5*5 = 125*5 = 625

OpenStudy (chosenmatt):

3 power of 2 plus 2 power of 4 @cwrw238

OpenStudy (chosenmatt):

simplify